OpenStudy (anonymous):

what is the derivative of f(x)=x(2x-1)^(1/2)

OpenStudy (anonymous):



OpenStudy (anonymous):

product derivation f(x)=g(x).h(x) f'(x)=g'(x).h(x) + g(x)h'(x)

OpenStudy (anonymous):

where g(x)=x and h(x)=(2x-1)^1/2

OpenStudy (anonymous):

product rule for this one \[(fg)'=f'g+g'f\] with \[f(x)=x,f'(x)=1,g(x)=\sqrt{2x-1},g'(x)=\frac{1}{\sqrt{2x-1}}\]

OpenStudy (anonymous):

anyway, wolfram alpha would help you with that kind of question

OpenStudy (anonymous):

aha i tried wolfram already but it was confusing. thanks guys